OpenStudy (anonymous):

Forty is 25% of what number?

OpenStudy (anonymous):

1.6 i think

OpenStudy (greencat):

\[40=\frac{25}{100}\times x\]

OpenStudy (greencat):

\[40\times \frac{100}{25}=x\]

OpenStudy (anonymous):

i think its 1.6

OpenStudy (greencat):

\[x=40\times4\]

OpenStudy (greencat):

x=160

OpenStudy (anonymous):

am i right

OpenStudy (greencat):

no

OpenStudy (anonymous):

so the answer is 160?

OpenStudy (anonymous):

This belongs in math

OpenStudy (anonymous):

@Great_And_Powerful no one helped so

OpenStudy (anonymous):

@GreenCat thanks

OpenStudy (greencat):

ur welcome

OpenStudy (greencat):

I think ur mistake was using 25 instead of 25%